Calculate the temperature of 30 g of di-atomic chlorine ( Cl2 )
gas that occupies a volume 800 L at a pressure of 10 kPa.
(atomic mass of chlorine is 35 u ).

2903.435 C
1975.996 C
3442.844 C
905.266 C
1255.705 C
1
Expert's answer
2019-07-05T14:15:18-0400

The classical ideal gas law states

"pV=\frac{m}{M}RT"

So, the temperature of di-atomic chlorine

"T=\frac{pVM}{Rm}"

"=\frac{10000\:\rm{Pa}\times 0.8\:\rm{m^3}\times 70\:\rm{g/mol}}{8.3\:\rm{J/(K\times mol)\times 30\:\rm{g}}}"

"=2248.996\:\rm{K}=1975.996\:^{\circ}\rm{C}"